OpenStudy (anonymous):

|2x-3|=11 ?? how do I do this one?

OpenStudy (anonymous):

4,7

OpenStudy (anonymous):

??? how do u work it out?

OpenStudy (anonymous):

2x-3=11 and 2x+3=11

OpenStudy (anonymous):

Well, as we know, since hgjweiryu told us ;0), there two answers. Absolute value tells us that if it -11 or +11 inside the absolute value then it is equal to 11. So Just see hgjweiryu second response.

OpenStudy (anonymous):

so it's 2x-3=11

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yes 2x-3=11 and 2x+3=11

OpenStudy (anonymous):

-(2x-3) = 11 and 2x-3 = 11. solve each of these for x.

OpenStudy (anonymous):

oh okay cause one - and one posttive rightt?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

hgjweiry is this correct?

OpenStudy (anonymous):

thankkkks yalll!